# Who’s important?—

### v2

@@ 1 Introduction
The abundance of online news media in the advent of the information era has led to a growing volume of daily data production [17], which poses significant challenges in efficiently identifying and understanding information concisely [1]. To promote efficacy and prevent information overloading, automated tasks such as Timeline Summarization (TLS) [4, 8, 6] becomes especially critical in this era [12, 23]. TLS is the generation of a summarized timeline of events, where multiple date-event pairs are sequentially listed to form a narrative for a particular topic. This process generally involves two essential requirements; (1) the events generated must be sequential and succinct (2) only relevant, fundamental and essential events should be included.
Several papers have attempted to improve the brevity of the first requirement and the identification of the second through the use of powerful Large Language Models (LLMs). These methods typically yield good results [22, 16, 6, 12, 23] due to the emergence of temporal understanding capabilities found within LLMs [24]. Nevertheless, the utilization of LLMs in TLS is still relatively underexplored [6, 16] as compared to works on temporal reasoning such as temporal question answering tasks [11].
@@
As such, we propose SUnSET: Synergistic Understanding of Stakeholder, Events and Time, for the task of automated TLS generation (Fig. 1). Unlike prior methods, SUnSET attempts to extract relevant information of various stakeholders mentioned within articles. Moreover, instead of summarizing articles immediately [12, 6], SUnSET attempts to extract multiple relevant events mentioned within an article to enable a better representation of occurrences documented by a single news article. The extracted stakeholder and event information will subsequently be combined with its estimated timeframe to generate a Stakeholder-Event-Time triplet (SET). This will be utilized in various clustering steps of the timeline generation process to eventually generate a singular summarized timeline. SUnSET was tested against classic TLS datasets [19, 20] and was able to improve existing baselines to emerge as the new state of the art.
@@ 2 Related Works
Recent TLS approaches can be broadly categorized into two main types: traditional methods and LLM-augmented methods; although variants of traditional clustering methods are still popular approaches when it comes to TLS task [5, 3], the rise of emergent capabilities in generative LLMs has led to increasing efforts to leverage these tools for TLS tasks, often with diverse objectives. Although LLM applications are more commonly associated with temporal reasoning, their usage in TLS has grown significantly over the past three years. For instance, Wu et al.’s paper highlighted the importance of temporal and causal relationships by incorporating LLMs into self-questioning strategies. On the other hand, works by Hu et al. and Qorib et al. introduced the notion of LLMs functioning as pseudo-oracles, emulating crowdsourced event clustering via pairwise querying during the clustering process.
Clustering methods in TLS are typically implemented in two distinct ways [5]. The more conventional approach involves grouping key dates to generate graph-based rankings [18] where subsequent summaries will be formed by selecting the top candidate sentences derived from generated ranked lists of important dates. Alternatively, the other method involves ‘event’ clustering [14], where article-level summaries generated once from each article are aggregated based on textual similarity. Increasingly, LLMs are integrated into these processes, especially to enhance summarization. For example, Hu et al. applied LLMs for article-level summarization prior to clustering, followed by a reclustering stage where the LLM validates cluster cohesion through pairwise comparisons of summary nodes. Meanwhile, Wu et al. explored iterative prompting techniques to survey context by progressively self-refining queries with LLMs to retrieve more relevant content. Building upon these efforts, our paper advances TLS clustering with LLM augmentation, where we will expand on stakeholder-based heuristics to assess event relevance.
@@ 3.1 SET Generation
The first step of SUnSET requires the generation of Stakeholder, Event and Time to form a triplet (SET). This is done via utilizing a LLM to extract events mentioned within an article and their estimated date of occurrence. Therefore, every article may contain multiple events: . After extracting the event and its estimated date (), we call the same LLM to extract relevant stakeholders mentioned within the article with a maximum of 5 stakeholders per event: . The term stakeholder used here strictly refers to an entity which is either a person or an organization. This results in a series of SET for a single article: . The relevant prompts used can be found in Appendix A.
Subsequently, Coreference Resolution will be done for all of the extracted stakeholders due to difference in naming such as utilizing a title or a position than a name (E.g. President of the United States) or differences in naming (E.g. POTUS v.s. President of America). A knowledge graph (KG) was built from the wikidata [21] database. The specifics of the KG building can be accessed in Appendix B.
@@ 3.2 Event Clustering
Reoccurrence |
Rare | Common | |
@@
The next stage of the workflow involves the event clustering process. Typically, this involves using content found within similarly dated articles to generate clusters, where each article forms an event node. Instead of using the first five sentences in the body of each article [5] or a pre-generated article summary [6, 12] as the input, SUnSET uses every single event () to generate a more accurate representation of an event’s impact.
@@
The penalty score () utilizes the coefficient of variation [2] to grasp the relative dispersion of stakeholder counts with respect to the sample mean across all topics. This value will be max-min normalized, and multiplied by the percentage of occurrence in the current topic. The boundary of and related proof will be in Appendix C. On the other hand, the reward score () uses a dampened hyperbolic tangent score with a factor of to limit the reward value from exploding, where the maximum possible value is capped at . The final relevance score thus multiplies the reward with the penalty score, before using a hyperparameter to tune the significance of the relevance score.
To test the efficacy of the aforementioned penalty score, we introduce another penalty scoring, which is an altered version of the inverse document frequency (IDF) found in BM25 [15]. Unlike Equation 2, Equation 5 directly uses the absence and presence of stakeholder counts across all documents to estimate the degree to penalize; this value however, fails to capture the actual extent of topic-based counts since it does not consider topic-based absence but document-based absence across the full document set. This may lead to edge cases being misrepresented: cases with common rarity and low reoccurrence (Irrelevant) may have same penalty output as those with both rare and high reoccurrence (Significant) since both cases provides high values. Contrary to using , the coefficient of variation in the original penalty score represents the spread of across all topics, which allows the subsequent penalization of compact counts while identifying low and high frequency topics in high-variation counts through percentage multiplications. The behaviour of the different penalty and reward scores will be further illustrated in Appendix D.
To increase the strictness of relevant events within the same cluster, we have also experimented with incorporating stakeholder Exact Matching (EM), where every 2 nodes requires at least N-matching unique stakeholders to have an edge. Therefore, the clustering process uses the encoded event summary from a General Text Embedding (GTE) Model [25] to obtain query-based cosine similarity scores combined with relevancy scores to generate the top 20 similar events for every node. These events have to satisfy the boolean EM condition to gain a weighted connection () with the query node (Eqn. 4). Due to the use of relevancy scores, our method does not need to incorporate resource-intensive pairwise LLM comparisons of events found in the same cluster [6, 12].
@@ 3.3 Timeline Generation
After obtaining multiple clusters, each cluster undergoes a re-ranking based on both their size and the existing relevance of their events’ stakeholders (Eqn. 7); instead of incorporating relevance scoring to aid in edge connection (Eqn. 4), the relevance score is applied to determine how significant each cluster is relative to both and . Events which are measured as significant will then be passed into TextRank [9] to identify important nodes within existing clusters. The final set of nodes will subsequently be used in the final Timeline Generation (TLG).
@@ 4 Experimental Setups
We used two renowned TLS datasets, Timeline17 (T17) [19] and Crisis [20]. T17 contains 19 timelines compiled from varying sources of online news sites, spanning 9 major topics from 2005-2013, each with 1-5 ground truth timelines. On the other hand, Crisis has 22 annotated timelines covering 4 critical crisis events, each containing 4-7 ground truth timelines.
@@
Adhering to previous work done [6, 12, 23], we incorporate a part of the Tilse framework [18] and use three main scores to analyse the performance of the TLS task. We use an Alignment-based ROUGE-1 F1-score (AR-1) to evaluate the semantic distance of unigram overlaps between generated timelines and the provided ground truth. We also used the corresponding bigram overlaps and scored the Alignment-based ROUGE-2 F1-score (AR-2). Lastly, the Date-F1 metric will score the similarity of the dates in the generated timeline compared to the referenced ground truth to understand the quality of major events picked up.
@@
Following prior papers [6, 12], the TLS task can be defined as follows: given a set of temporally labelled news articles that is related to a broad topic (), as well as the expected number of dates and the number of sentences to include in each date, a single timeline summarizing the topic should be generated. This generated timeline will subsequently be compared with the ground truth provided in the datasets. The baselines from the selected benchmarks are run in their publicly available GitHub repository.
We will release our full code publicly on Github once the paper is accepted into the conference.
@@ 5 Experimental Results
Table 2 showcases SUnSET as compared to the previously SOTA baselines. We selected the strongest model for each method for the comparisons. We also included Llama2-13B in Table 2 since Llama2 was the original model used in LLM-TLS’s paper [6] and it occasionally performs better than Qwen2.5-72B.
@@ 5.1 Relevancy in Event Clustering
We further investigated the impact of relevance used in the event clustering process by ablating all stakeholder-related information and relevancy scores, retaining only the cosine similarity scores of the events. From the results observed in Table 3, the use of relevance scoring and stakeholder information causes a leap in performance on Date-F1 scores by for Crisis and for T17. This indicates that our method for the identification of stakeholders plays an integral part in ranking event importance. Moreover, even without implementing , it is observed that utilizing LLMs to generate event sets in SET is already more effective than existing methods which typically uses summarization of [12, 6] or a self-questioning rewriter [23]. Even without pairwise matching (“w/o ”), using only the cosine similarity of managed to beat the benchmark set by LLM-TLS and CHRONOS (Tab. 2). This reveals that extraction of multiple mentioned events within a single article is vital for TLS tasks; this is intuitive, as news articles often reference related developments that connect to the central event being reported. Significant occurrences may receive limited coverage initially, until a subsequent development—part of an ongoing sequence—draws renewed attention from multiple news outlets.
@@ 5.5 Time Comparison
Lastly, we measured the GPU hours required for running the experiments, and realised the high efficiency for using SUnSET as compared to previous SOTA methods. The benefits mainly comes from the lack of additional LLM prompting beyond SET creation, where CHRONOS utilizes questioning rounds and LLM-TLS prompts once for every pairwise comparison required. Therefore, SUnSET is able to bypass such time-consuming prompt-questioning methods due to the high efficacy of utilizing stakeholder information for distilling events into a compact representation.
@@ 6 Conclusions
Our study advances TLS through the integration of SET and stakeholder ranking mechanisms, though certain limitations remain. Within our domain of control, we applied feasible refinement techniques such as prompt engineering and performed rigorous reviews to minimize inconsistencies. However, as with prior work, we were unable to fully mitigate hallucinations or factual inaccuracies that may arise during event extraction when utilizing LLMs due to their inherent generation behavior. Additionally, our experiments were constrained to two representative LLMs, one open-source and one closed-source model, which limits broader generalization across models. Finally, in line with previous studies, we evaluated SUnSET using publicly available research datasets, without extending analysis to real-time news data with more dynamic, evolving topics. We identify this as a promising direction for future exploration.

### v3

@@ Abstract
Who’s important?—
: Synergistic Understanding of Stakeholder,
Events and Time for Timeline Generation
Tiviatis Sim1,2, Yang Kaiwen3, Shen Xin3, Kenji Kawaguchi1 1National University of Singapore, 2A*STAR Institute of High Performance Computing, 3Celia Research Team, Norbert Wiener Research Center, Huawei International tiviatis@u.nus.edu, kaiwen_yang@alumni.brown.edu, sxstar@zju.edu.cn, kenji@comp.nus.edu.sg,
@@ 1 Introduction
The abundance of online news media in the advent of the information era has led to a growing volume of daily data production (Statista, ), which poses significant challenges in efficiently identifying and understanding information concisely (Arnold et al., 2023). To promote efficacy and prevent information overloading, automated tasks such as Timeline Summarization (TLS) (Chieu and Lee, 2004; Li et al., 2021; Hu et al., 2024) becomes especially critical in this era (Qorib et al., 2025; Wu et al., 2025). TLS is the generation of a summarized timeline of events, where multiple date-event pairs are sequentially listed to form a narrative for a particular topic. This process generally involves two essential requirements; (1) the events generated must be sequential and succinct (2) only relevant, fundamental and essential events should be included.
Several papers have attempted to improve the brevity of the first requirement and the identification of the second through the use of powerful Large Language Models (LLMs). These methods typically yield good results (Wang et al., 2023; Sojitra et al., 2024a; Hu et al., 2024; Qorib et al., 2025; Wu et al., 2025) due to the emergence of temporal understanding capabilities found within LLMs (Xiong et al., 2024). Nevertheless, the utilization of LLMs in TLS is still relatively underexplored (Hu et al., 2024; Sojitra et al., 2024a) as compared to works on temporal reasoning such as temporal question answering tasks (Piryani et al., 2025).
@@
As such, we propose SUnSET: Synergistic Understanding of Stakeholder, Events and Time, for the task of automated TLS generation (Fig. 1). Unlike prior methods, SUnSET attempts to extract relevant information of various stakeholders mentioned within articles. Moreover, instead of summarizing articles immediately (Qorib et al., 2025; Hu et al., 2024; Sojitra et al., 2024b), SUnSET attempts to extract multiple relevant events mentioned within an article to enable a better representation of occurrences documented by a single news article. The extracted stakeholder and event information will subsequently be combined with its estimated timeframe to generate a Stakeholder-Event-Time triplet (SET). This will be utilized in various clustering steps of the timeline generation process to eventually generate a singular summarized timeline. SUnSET was tested against classic TLS datasets (Tran et al., 2013, 2015) and was able to improve existing baselines to emerge as the new state of the art.
@@ 2 Related Works
Recent TLS approaches can be broadly categorized into two main types: traditional methods and LLM-augmented methods; although variants of traditional clustering methods are still popular approaches when it comes to TLS task (Gholipour Ghalandari and Ifrim, 2020; Chen et al., 2023), the rise of emergent capabilities in generative LLMs has led to increasing efforts to leverage these tools for TLS tasks, often with diverse objectives. Although LLM applications are more commonly associated with temporal reasoning, their usage in TLS has grown significantly over the past three years. For instance, Wu et al.’s paper highlighted the importance of temporal and causal relationships by incorporating LLMs into self-questioning strategies. On the other hand, works by Hu et al. and Qorib et al. introduced the notion of LLMs functioning as pseudo-oracles, emulating crowdsourced event clustering via pairwise querying during the clustering process.
Clustering methods in TLS are typically implemented in two distinct ways (Gholipour Ghalandari and Ifrim, 2020). The more conventional approach involves grouping key dates to generate graph-based rankings (Steen and Markert, 2019) where subsequent summaries will be formed by selecting the top candidate sentences derived from generated ranked lists of important dates. Alternatively, the other method involves ‘event’ clustering (Ribeiro et al., 2017), where article-level summaries generated once from each article are aggregated based on textual similarity. Increasingly, LLMs are integrated into these processes, especially to enhance summarization. For example, Hu et al. applied LLMs for article-level summarization prior to clustering, followed by a reclustering stage where the LLM validates cluster cohesion through pairwise comparisons of summary nodes. Meanwhile, Wu et al. explored iterative prompting techniques to survey context by progressively self-refining queries with LLMs to retrieve more relevant content. Building upon these efforts, our paper advances TLS clustering with LLM augmentation, where we will expand on stakeholder-based heuristics to assess event relevance.
@@ 3.1 SET Generation
The first step of SUnSET requires the generation of Stakeholder, Event and Time to form a triplet (SET). This is done via utilizing a LLM to extract events mentioned within an article and their estimated date of occurrence. Therefore, every article may contain multiple events: . After extracting the event and its estimated date (), we call the same LLM to extract relevant stakeholders mentioned within the article with a maximum of 5 stakeholders per event: . The term stakeholder used here strictly refers to an entity which is either a person or an organization. This results in a series of SET for a single article: . The relevant prompts used can be found in Appendix A. To minimize disruptions from hallucination, we used heuristic to ensure the generated date and event exists within the article and removed illogical dates. For events with lack of reference dates due to the lack of information within the article, the event uses the date of publication.
Subsequently, Coreference Resolution will be done for all of the extracted stakeholders due to difference in naming such as utilizing a title or a position than a name (E.g. President of the United States) or differences in naming (E.g. POTUS v.s. President of America). A knowledge graph (KG) was built from the wikidata (Vrandečić and Krötzsch, 2014) database. This allows us to capture entities by origin-language forms and transliterations common within non-English dominant regions such as Semitic languages. The specifics of the KG building can be accessed in Appendix B.
@@ 3.2 Event Clustering
| Reoccurrence | Rare | Common | |
@@
The next stage of the workflow involves the event clustering process. Typically, this involves using content found within similarly dated articles to generate clusters, where each article forms an event node. Instead of using the first five sentences in the body of each article (Gholipour Ghalandari and Ifrim, 2020) or a pre-generated article summary (Hu et al., 2024; Qorib et al., 2025) as the input, SUnSET uses every single event () to generate a more accurate representation of an event’s impact. This clustering process automatically reduces error brought forward by hallucination not caught in the previous step since hallucinations generate small event clusters which would be neglected due to their small event size.
@@
The penalty score () utilizes the coefficient of variation (Brown, 1998) to grasp the relative dispersion of stakeholder counts with respect to the sample mean across all topics. This value will be max-min normalized, and multiplied by the percentage of occurrence in the current topic. The boundary of and related proof will be in Appendix C. On the other hand, the reward score () uses a dampened hyperbolic tangent score with a factor of to limit the reward value from exploding, where the maximum possible value is capped at . The final relevance score thus multiplies the reward with the penalty score, before using a hyperparameter to tune the significance of the relevance score.
To test the efficacy of the aforementioned penalty score, we introduce another penalty scoring, which is an altered version of the inverse document frequency (IDF) found in BM25 (Robertson and Zaragoza, 2009). Unlike Equation 2, Equation 5 directly uses the absence and presence of stakeholder counts across all documents to estimate the degree to penalize; this value however, fails to capture the actual extent of topic-based counts since it does not consider topic-based absence but document-based absence across the full document set. This may lead to edge cases being misrepresented: cases with common rarity and low reoccurrence (Irrelevant) may have same penalty output as those with both rare and high reoccurrence (Significant) since both cases provides high values. Contrary to using , the coefficient of variation in the original penalty score represents the spread of across all topics, which allows the subsequent penalization of compact counts while identifying low and high frequency topics in high-variation counts through percentage multiplications. The behaviour of the different penalty and reward scores will be further illustrated in Appendix D.
To increase the strictness of relevant events within the same cluster, we have also experimented with incorporating stakeholder Exact Matching (EM), where every 2 nodes requires at least N-matching unique stakeholders to have an edge. Therefore, the clustering process uses the encoded event summary from a General Text Embedding (GTE) Model (Zhang et al., 2024) to obtain query-based cosine similarity scores combined with relevancy scores to generate the top 20 similar events for every node. These events have to satisfy the boolean EM condition to gain a weighted connection () with the query node (Eqn. 4). Due to the use of relevancy scores, our method does not need to incorporate resource-intensive pairwise LLM comparisons of events found in the same cluster (Hu et al., 2024; Qorib et al., 2025).
@@ 3.3 Timeline Generation
After obtaining multiple clusters, each cluster undergoes a re-ranking based on both their size and the existing relevance of their events’ stakeholders (Eqn. 7); instead of incorporating relevance scoring to aid in edge connection (Eqn. 4), the relevance score is applied to determine how significant each cluster is relative to both and . Events which are measured as significant will then be passed into TextRank (Mihalcea and Tarau, 2004) to identify important nodes within existing clusters. The final set of nodes will subsequently be used in the final Timeline Generation (TLG).
Henceforth, variations of overlapping events over different articles will not be a pressing concern since the clustering process is able to identify these variations based both the score from similar texts and their scores from repeated stakeholders to create a tighter cluster. In addition, TextRank, which causes nodes which are more different but within the same cluster to be used, would balance the variations in duplicates within their own cluster to better represent the events which are same with differing narrations. The penalty represented in also helps prevent similar but different events from falling into wrong cluster.
@@ 4 Experimental Setups
We used two renowned TLS datasets, Timeline17 (T17) (Tran et al., 2013) and Crisis (Tran et al., 2015). T17 contains 19 timelines compiled from varying sources of online news sites, spanning 9 major topics from 2005-2013, each with 1-5 ground truth timelines. On the other hand, Crisis has 22 annotated timelines covering 4 critical crisis events, each containing 4-7 ground truth timelines.
@@
For experiments, we adopted Qwen2.5-72B-Instruct (Qwen et al., 2025), and GPT-4 omni (OpenAI, 2024). We used GTE-Modernbert-Base for encoding. For the main experimental results, we used the best model, Qwen2.5-72B-Instruct to compare across the different benchmarks. All model deployment were done with VLLM (Kwon et al., 2023) with 8 H200 GPUs. We used batch sizes of 256 for batch encoding, and limit each event summary to a maximum of 4096 characters.
@@
Adhering to previous work done (Hu et al., 2024; Qorib et al., 2025; Wu et al., 2025), we incorporate a part of the Tilse framework (Steen and Markert, 2019) and use three main scores to analyse the performance of the TLS task. We use an Alignment-based ROUGE-1 F1-score (AR-1) to evaluate the semantic distance of unigram overlaps between generated timelines and the provided ground truth. We also used the corresponding bigram overlaps and scored the Alignment-based ROUGE-2 F1-score (AR-2). Lastly, the Date-F1 metric will score the similarity of the dates in the generated timeline compared to the referenced ground truth to understand the quality of major events picked up.
@@
Following prior papers (Hu et al., 2024; Qorib et al., 2025), the TLS task can be defined as follows: given a set of temporally labelled news articles that is related to a broad topic (), as well as the expected number of dates and the number of sentences to include in each date, a single timeline summarizing the topic should be generated. This generated timeline will subsequently be compared with the ground truth provided in the datasets. The baselines from the selected benchmarks are run in their publicly available Github repository. We have also published SUnSET publicly on Github for reproduction.
@@ 5 Experimental Results
Table 2 showcases SUnSET as compared to the previously SOTA baselines. We selected the strongest model for each method for the comparisons. We also included Llama2-13B in Table 2 since Llama2 was the original model used in LLM-TLS’s paper (Hu et al., 2024) and it occasionally performs better than Qwen2.5-72B.
@@ 5.1 Relevancy in Event Clustering
We further investigated the impact of relevance used in the event clustering process by ablating all stakeholder-related information and relevancy scores, retaining only the cosine similarity scores of the events. From the results observed in Table 3, the use of relevance scoring and stakeholder information causes a leap in performance on Date-F1 scores by for Crisis and for T17. This indicates that our method for the identification of stakeholders plays an integral part in ranking event importance.
Moreover, even without implementing , by comparing Table 2’s LLM-TLS score and Table 3’s “w/o ”, it illustrates the significance of the event component (E). It is observed that utilizing LLMs to generate event sets in SET is already more effective than existing methods which typically uses summarization of (Qorib et al., 2025; Hu et al., 2024) or a self-questioning rewriter (Wu et al., 2025). Even without pairwise matching (“w/o ”), using only the cosine similarity of managed to beat the benchmark set by LLM-TLS and CHRONOS (Tab. 2). This reveals that extraction of multiple mentioned events within a single article is vital for TLS tasks; this is intuitive, as news articles often reference related developments that connect to the central event being reported. Significant occurrences may receive limited coverage initially, until a subsequent development—part of an ongoing sequence—draws renewed attention from multiple news outlets. Therefore, since the time component in SET is always a necessary input due to the nature of this task, we have shown that all components in SET are important for the performance of timeline generation.
@@ 5.5 Time Comparison
Lastly, we measured the time required for running the experiments, and realised the high efficiency for using SUnSET as compared to previous SOTA methods. The benefits mainly comes from the lack of additional LLM prompting beyond SET creation, where CHRONOS utilizes questioning rounds and LLM-TLS prompts once for every pairwise comparison required. Therefore, SUnSET is able to bypass such time-consuming prompt-questioning methods due to the high efficacy of utilizing stakeholder information for distilling events into a compact representation.
@@ 6 Conclusions
Our study advances TLS through the integration of SET and stakeholder ranking mechanisms, though certain limitations remain. Within our domain of control, we applied feasible refinement techniques such as prompt engineering and performed rigorous reviews to minimize inconsistencies. However, as with prior work, we were unable to fully mitigate hallucinations or factual inaccuracies that may arise during event extraction when utilizing LLMs due to their inherent generation behavior. Finally, in line with previous studies, we evaluated SUnSET using publicly available research datasets, without extending analysis to real-time news data with more dynamic, evolving topics. We identify this as a promising direction for future exploration.